OpenStudy (jjuden):

Which of the following methods would be the easiest to use to solve 2x2 + 4x – 3 = 0?

OpenStudy (jjuden):

A factoring B isolating the x2 term and finding the square root of both sides C using the quadratic formula D all three methods would be easy and effective

OpenStudy (accessdenied):

What have you tried with this problem so far? I'd just go through and try each step. The quadratic formula always works with the right information, so we should just try (A) and (B).

OpenStudy (johnweldon1993):

Since the problem won't deal with whole numbers...it would be easiest to use the quadratic formula

OpenStudy (jjuden):

ok thank you

OpenStudy (johnweldon1993):

That is just my opinion however...it WOULD be best to go through each option and try each one out to see which one works easiest for you

OpenStudy (anonymous):

me, i am a big fan of completing the square \[2x^2 + 4x – 3 = 0\\ 2x^2+4x=3\\ x^2+2x=\frac{3}{2}\\ (x+1)^2=\frac{3}{2}+1=\frac{5}{2}\] etc